Explain with reasoning.
You do not have the required permissions to view the files attached to this post.
Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.
See classesLearn live with an expert or move at your own pace. Every option includes the complete TTP study system.
Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.
with Josh Braslow, 20+ years in test prep
Self-paced EA prep. Study on your schedule.
with
Complete access from day one. Study on your schedule.
DanaJ, you did a calculation error. The number of favorable cases as per your explanation is 6 not 5.DanaJ wrote:I'd go with 5/32 and here's my reasoning:
There are favorable and unfavorable cases to each of the five tosses, and since each case has a probability of 1/2, you'll always have a total probability of 1/2*1/2*1/2*1/2*1/2 = 1/32 for your event. The distribution is as follows:
H H H T T
T H H H T
T T H H H
T H H H H
H H H H T
H H H H H
Since each of the cases above is obtained by having a suite of 5 favorable tosses, then the total probability will be 5*1/32 = 5/32.
IMO Bcramya wrote:I would go with A
3 heads or 4 heads or 5 heads
Pick 5 spots there are three ways to get 3 heads 9in consecutive tosses and 2 tails
1/32+1/32+1/32 = 3/32
For 4 consecutive heads there are 2 ways to get it in 5 tosses
2/32
5 consecutive heads there is only one choice = 1/32
3/32+2/32+1/32
= 6/32
= 3/16
Yes Cramya! Right!cramya wrote:Piyush,
HTHHH
HHHTH
These wont be consecutive tosses, correct?
Regards,
CR